METHANOL IN TRANSLUCENT INTERSTELLAR MOLECULAR CLOUDS

\underline{B.E. TURNER}, NRAO, 520 Edgemont Rd., Charlottesville, VA 22903, USA.

We have detected several low-J lines of E- and A-type methanol in 17 translucent interstellar clouds whose physical conditions and chemistry have recently been well studied. The mean fractional abundance is 1(-8), similar to the abundance of 3(-9) found in cold dark clouds. Gas-phase formation occurs via CH_{3}^{+} + H_{2}O → CH_{3}OH_{2}^{+} + hv. We have verified the predictions of this simple process by using the full Standard Model of interstellar chemistry (over 3000 reactions). While most interstellar molecules are formed by gas-phase processes, these models predict CH_{3}OH abundances 4 orders of magnitude less than the observed abundances. We have examined grain surface chemistry in which accreted CO hydrogenates to CH_{3}OH on the surface under the action of UV or cosmic rays, then desorbs principally by photodesorption. These processes can easily explain the observed CH_{3}OH abundances, and in fact imply lower desorption efficiencies than usually adopted.